The Central Bank has collected the following monetary statistics for the year 2015: Notes and coins Kshs.800 billion, Demand deposits Kshs.1,600.billion, Fixed Deposits Kshs.2,400 billion, Deposits with NBFIs Kshs.3,000 billions; Foreign currency deposits of citizens and foreigners Kshs.2,000 billions, 91 days Treasury Bills 3,000 billions, 182 Treasury Bills 4,000 billions, 300 days Treasury Bills 4,000 billions. Bank deposits with Central Bank 1,200 Billions. Page 1 of 2

media%2Fab8%2Fab8fc6e7-abfb-43f6-a99d-d9
Answer #1

Solution:

M1, M2, M3, M3X, M3XT these are basically the measures to define monetary aggregates of the country as per their liquidity.

M1 = M0 + Demand deposits including sight deposits of trade and industry, commercial banks and postal accounts

where M0 is currency in circulation basically notes and coins

M1= 800 + 1600 = 2400 billion

M2 = M1 + Time deposits like FD

M2= 2400 + 2400 = 4800 billion

M3= M2 + Non bank financial institution (NBFI)

M3 = 4800 + 3000 = 7800 billion

M3X is also called as broad money supply = M3 + Residents foreign currency deposit

M3X= 7800 + 2000 = 9800 billion

M3XT also called as broad liquidity measure = M3X + T bills

M3XT = 9800 + 11000 = 20800 billion

Solution 2:

r0= 5% and rT=10%

m 1= 1+(C/D)/[rr+(ER/D)+(C/D)]

= 1+ 800/1600/[(0.05+(1200/1600) + (800/1600)]

1+ (0.5/[0.05+0.75+0.5)

=1+ 0.5/1.3= 1+0.384=1.384

m2=1+(C/D)+(T/D)+(MMF/D)/[rr+(ER/D)+(C/D)]

1+0.5+2400/1600+(3000/1600)/ 0.1+0.75+0.5

(1+0.5+1.5+1.875)/1.35

4.875/1.35=3.611

Money multiplier is basically the ratio of broad money and reserve money (M0). Here in this case the broad money is M2

Money multiplier= M2/M0= 4800/2400 = 2